Two classes, one with 50 students and the other with 30, take the same exam. The combined average of both classes is 84.5. If the larger class averages 80, what is the average of the smaller class?
Answer by checkley71(8403)

80*84.5=6760 OR 50*80+30X=6760 OR 4000+30X=6760 OR 30X=2760 OR X=92 THE AVERAGE GRADES FOR THE 30 STUDENT CLASS.